OpenStudy (anonymous):

How do I do this and what is the answer? If g(x) = x2 + 3, find g(4). A) 11 B) 19 C) 16 D) 8

OpenStudy (anonymous):

g(x) = x^2 + 3, find g(4) So, 4 = 4^2 + 3. Can you simplify from there?

OpenStudy (anonymous):

not really. could you do it for me @EllenJaz17 ?

OpenStudy (anonymous):

g(4)=4^2+3=16+3=19

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

could you help me with a few more problems @EllenJaz17 ?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Ok Thanks! This is the first one: If g(x) = 2(x - 4), find the value of x if g(x) = 20. A) 32 B) 12 C) 14 D) 10

OpenStudy (anonymous):

20 = 2x-8 20+8=2x 28=2x 28/2=x 14=x

OpenStudy (anonymous):

@jut21

OpenStudy (anonymous):

Can you give me a medal please lol

OpenStudy (anonymous):

thanks! and yah sorry I forgot about that

OpenStudy (anonymous):

I have more problems that I need help with. is that ok?

OpenStudy (anonymous):

@EllenJaz17

OpenStudy (anonymous):

yeah @jut21

OpenStudy (anonymous):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie. A) c(x) = 2.00x + 0.50 B) c(x) = 2.00 + 0.50x C) c(x) = 2.50x D) c(x) = (2.00 + 0.50)x

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s. Which statement best describes the meaning of f(150)? A) This is the average number of days the house stayed on the market before being sold for $150,000. B) The house sold on the market for $150,000 and stayed on the market for an average of 150 days before being sold. C) The house sold for $150,000. D) The house stayed on the market for an average of 150 days before being sold.

OpenStudy (anonymous):

@EllenJaz17

OpenStudy (anonymous):

My best guess is B because it includes both variables in the question @jut21

OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. A) 2, 3, 5, 7, 9 B) 2, 3, 4, 7, 11 C) 2, 3, 5, 9, 11 D) 2, 3, 5, 8, 13

OpenStudy (anonymous):

@EllenJaz17

OpenStudy (anonymous):

Sorry Im doing my own hw as well lol @jut21